# Diff-BBO: Diffusion-Based Inverse Modeling for Black-Box Optimization

### v1

@@ Abstract
  1. Black-box optimization (BBO) aims to optimize an objective function by iteratively querying a black-box oracle. This process demands sample-efficient optimization due to the high computational cost of function evaluations. While prior studies focus on forward approaches to learn surrogates for the unknown objective function, they struggle with high-dimensional inputs where valid inputs form a small subspace (e.g., valid protein sequences), which is common in real-world tasks. Recently, diffusion models have demonstrated impressive capability in learning the high-dimensional data manifold. They have shown promising performance in black-box optimization tasks but only in offline settings. In this work, we propose diffusion-based inverse modeling for black-box optimization (Diff-BBO), the first inverse approach leveraging diffusion models for online BBO problem. Diff-BBO distinguishes itself from forward approaches through the design of acquisition function. Instead of proposing candidates in the design space, Diff-BBO employs a novel acquisition function Uncertainty-aware Exploration (UaE) to propose objective function values, which leverages the uncertainty of a conditional diffusion model to generate samples in the design space. Theoretically, we prove that using UaE leads to optimal optimization outcomes. Empirically, we redesign experiments on the Design-Bench benchmark for online settings and show that Diff-BBO achieves state-of-the-art performance.
@@ 1 Introduction
Practical problems in science and engineering often involve optimizing a black-box objective function that is expensive to evaluate, such as neural network architecture design (Zoph and Le, 2016), robotics (Tesch et al., 2013), and molecular design (Sanchez-Lengeling and Aspuru-Guzik, 2018). How to achieve a near-optimal solution while minimizing function evaluations is thus a major challenge in black-box optimization (BBO). To improve sample efficiency, sequential optimization methods iteratively generate new samples in the design space conditioning on existing data and query the black-box function for evaluation. One of the most prevailing approaches is Bayesian optimization (BO) (Kushner, 1964; Mockus, 1974). It is a class of forward methods that learns a surrogate model for the objective function with Gaussian processes (GPs), and applies an acquisition function based on the surrogate to select new data samples (Frazier, 2018).
@@
Despite its efficacy, the curse of dimensionality limits the applicability of BO to high-dimensional problems in terms of both statistical and computational complexities (Malu et al., 2021). Additionally, these approaches can struggle with high-dimensional inputs where valid inputs constitute a small subspace (e.g., valid protein sequences or molecular structures). Such optimization problems become exceptionally challenging, as the optimizer must navigate and avoid out-of-distribution and invalid inputs (Kumar and Levine, 2020). Recently, a novel set of methods, termed inverse approaches, have been proposed to address this issue. These methods (Kumar and Levine, 2020; Krishnamoorthy et al., 2023; Kim et al., 2024) break the traditional paradigm by learning an inverse mapping from objective space back to the input space (a.k.a., the black-box function’s design domain). Leveraging state-of-the-art generative models, such as diffusion models (Song et al., 2020), these approaches effectively capture data distributions in high-dimensional input space and facilitate optimization within the data manifold (Kong et al., 2024). Besides, diffusion models naturally provide uncertainty estimates through the probabilistic nature of the diffusion process, which can be further utilized to design informative exploration strategies to propose better candidate solutions for optimization problems. However, existing inverse methods that leverage diffusion models have only been studied in offline optimization settings, assuming access to a fixed pre-collected dataset.
In this paper, we propose Diff-BBO, an inverse approach for online black-box optimization. Diff-BBO places a distribution on the design space and represents it with a conditional diffusion model. It leverages the uncertainty of the diffusion model to query the oracle function efficiently. Compared to the existing algorithms, it enjoys scalability in solving high-dimensional Black-box Optimization (BBO) problems, sample efficiency in online settings, as well as the capability of quantifying uncertainty in inverse modeling. Our approach consists of a novel acquisition function design through the uncertainty quantification (UQ) of conditional diffusion model, which proposes the desired objective function values used to strategically sample the design space. We summarize our main contributions as follows:
  • We present Diff-BBO, an inverse modeling approach for efficient online black-box optimization (BBO) leveraging uncertainty of conditional diffusion models.
@@
  • We provide an uncertainty decomposition into epistemic uncertainty and aletoric uncertainty for conditional diffusion models. We rigorously analyze how uncertainty propagates throughout the denoising process of conditional diffusion model.
@@
  • We design a novel acquisition function based on the uncertainty for BBO. Theoretically, we prove that the balance between targeting higher objective values and minimizing epistemic uncertainty lead to optimal optimization outcomes.
@@
  • We demonstrate that Diff-BBO achieves state-of-the-art performance with superior sample efficiency on Design-Bench and molecular discovery task in the online BBO setting.
@@ 2 Related Work
Black-box Optimization.
While recent studies aim to solve offline Black-box Optimization (BBO) using a pre-collected dataset (Li et al., 2024; Krishnamoorthy et al., 2023; Fu and Levine, 2021) without querying the oracle function, prior works in BBO have largely focused on the online setting where a model can iteratively query the function during training (Turner et al., 2021; Zhang et al., 2021; Hebbal et al., 2019; Mockus, 1974). In online BBO, most existing algorithms belong to forward methods, including BO (Wu et al., 2023) and bandit algorithms (Agrawal and Goyal, 2012; Karbasi et al., 2023). Forward methods build a surrogate model to approximate the black-box function and optimize sequentially. Song et al.; Zhang et al. proposed Likelihood-free BO uses likelihood-free inference to extend BO to a broader class of models and utilities. It directly models the acquisition function without separately performing inference with a surrogate model. However, there is a risk where the acquisition function is over-confident. In light of the recent progress in inverse approaches (Krishnamoorthy et al., 2023; Kong et al., 2024), our work focuses on solving the online BBO by introducing a sample-efficient inverse modeling method using uncertainty-aware conditional diffusion models.
@@
As an emerging class of generative models with strong expressiveness, diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2020) have been successfully deployed across various domains including image generation (Rombach et al., 2022), reinforcement learning (Wang et al., 2022), robotics (Chi et al., 2023), etc. Notably, through the formulation of stochastic differential equations (SDEs), (Song et al., 2020) provides a unified continuous-time score-based framework for distinctive classes of diffusion models. To steer the generation towards high-quality samples of desired properties, guiding the backward data-generation process with encoded task-specific information plays an important role. Hence, different types of guidance are studied in prior works (Bansal et al., 2023; Nichol et al., 2021), including classifier guidance (Dhariwal and Nichol, 2021) where the classifier is trained externally, and classifier-free guidance (Ho and Salimans, 2022), in which the classifier is implicitly specified. In this work, we focus on classifier-free guidance to provide feasible uncertainty quantification for conditional diffusion models.
@@
Reliable uncertainty quantification (UQ) often relies on probabilistic modeling, with Bayesian approximation and ensemble learning being two popular types of approaches. Bayesian Neural Networks (BNNs) (MacKay, 1992; Neal, 2012; Kendall and Gal, 2017; Zhang et al., 2018) operate by employing variational inference (VI) to sample model weights from a tractable distribution and estimate uncertainty through sample variance. When training large-scale models, Monte Carlo (MC) dropout (Srivastava et al., 2014) offers a cost-effective alternative by approximating BNNs during inference (Gal and Ghahramani, 2016). On the other hand, deep ensembles (Lakshminarayanan et al., 2017) train multiple NNs with different initial weights to gauge uncertainty via model variance, which also faces scalability issues as network size increases. To address this issue, recent efforts have been made to incorporate ensembling techniques with generative models to separate uncertainty into aleatoric and epistemic components (Valdenegro-Toro and Mori, 2022; Ekmekci and Cetin, 2023). To further improve the scalability of deep ensembles, (Chan et al., 2024) proposed hyper-diffusion to quantify the uncertainty with a single diffusion model. In comparison, we take one step further by utilizing the quantified uncertainty of conditional diffusion models to solve the black-box optimization problem as a downstream task.
3 Diff-BBO
In this section, we present Diffusion-based Inverse Modeling for Black-Box Optimization (Diff-BBO). We first provide a probabilistic formulation of the online BBO problem along with our algorithm, followed by details of training conditional diffusion model in Diff-BBO to obtain uncertainty.
Figure 1: Forward modeling vs inverse modeling for black-box optimization. (Top) Forward modeling approach using certain surrogate models (e.g., GPs) for forward modeling and acquisition functions (e.g., UCB, PI, and EI) to select $\mathbf{x}$. (Bottom) Our inverse modeling approach using generative model (e.g., diffusion model) for inverse modeling and acquisition function (e.g., UaE) to select $y$.
3.1 The Diff-BBO Algorithm
@@
Let $f:\mathcal{X}\rightarrow\mathbb{R}$ denote the unknown ground-truth black-box function that evaluates the quality of any data point $\bm{x}$, with $\mathcal{X}\subseteq\mathbb{R}^{d}$. Our goal is to find the optimal point $\bm{x}^{*}$ that maximizes $f$:
$\bm{x}^{*}\in\operatorname*{argmax}\nolimits_{\bm{x}\in\mathcal{X}}f(\mathbf{\bm{x}}).$
@@
We are interested in the online BBO setting in which $f$ is expensive to evaluate and the number of evaluations is limited. In particular, we consider batch online BBO. With a fixed query budget of $Q$ and batch size $N$, we iteratively query $f$ with $N$ new inputs in each bath, and update the model based on observed outputs within $K$ iterations. A key concept in online BBO is the acquisition function, which guides the selection of new query points by balancing exploration and exploitation. This function aims to efficiently identify high-performing inputs, thereby efficiently solving the online optimization problem.
In our inverse modeling approach, we model the conditional distribution of $p(\bm{x}|y,\mathcal{D})$ with training data $\mathcal{D}$. The function value $y$ to condition on is proposed by an acquisition function, which quantifies the quality of the generated $\bm{x}$. The objective of the above optimization becomes:
$\max_{y_{k}\in\mathbb{R}}\sum_{k=1}^{K}f(\bm{x}_{k}),~{}~{}\bm{x}_{k}\sim p_{\theta}(\cdot\mid y_{k},\mathcal{D}),~{}~{}\theta\in\Theta.$ (1)
To solve the above optimization problem, we introduce Diff-BBO in Algorithm 1. At each iteration $k$, we train a conditional diffusion model and compute the optimal $y_{k}^{*}$ with the designed acquisition function. In practice, $y$ is selected from a constructed candidate set $\mathcal{Y}$ based on the acquisition function scores $\alpha(y)$. The range of $\mathcal{Y}$ is determined by $w\cdot\phi_{k}$, where $w$ is a positive scalar and $\phi_{k}$ is the maximum function values being queried in the current training dataset $\mathcal{D}$. Conditioning on $y_{k}^{*}$, we generate $N$ samples $\{\mathbf{x}_{j}\}_{j=1}^{N}$, where $\bm{x}_{j}\sim p_{\theta}(\mathbf{x}|y_{k}^{*},\mathcal{D})$. By querying the black-box oracle to evaluate each $\bm{x}_{j}$, we obtain the best possible reconstructed value $\phi_{k}$ for the current iteration, and append all queried data pairs $\{\mathbf{x}_{j},f(\mathbf{x}_{j})\}_{j=1}^{N}$ to the training dataset $\mathcal{D}$.
@@
Figure 1 demonstrates a detailed comparison of prior forward methods to solve the BBO problem with our proposed inverse modeling approach. Forward methods aim to learn the surrogate model $p(y|\bm{x},\mathcal{D})$ for the unknown objective $f$ by utilizing models such as GPs. However, these methods often encounter scalability issues and typically rely on heuristic approaches to generate new candidate solutions. In contrast, our approach leverages the power of diffusion model to represent $p(\bm{x}|y,\mathcal{D})$, allowing it to provide high-quality candidate solutions in high-dimensional problem domains and to condition on arbitrary function value. Besides, diffusion models naturally provide uncertainty estimates, which are further utilized in our design of acquisition function to solve the BBO problem.

### v3

# Diffusion-BBO: Diffusion-Based Inverse Modeling for Online Black-Box Optimization

@@ Abstract
  1. Online black-box optimization (BBO) aims to optimize an objective function by iteratively querying a black-box oracle in a sample-efficient way. While prior studies focus on forward approaches such as Gaussian Processes (GPs) to learn a surrogate model for the unknown objective function, they struggle with steering clear of out-of-distribution and invalid designs in scientific discovery tasks. Recently, inverse modeling approaches that map the objective space to the design space with conditional diffusion models have demonstrated impressive capability in learning the data manifold. However, these approaches proceed in an offline fashion with pre-collected data. How to design inverse approaches for online BBO to actively query new data and improve the sample efficiency remains an open question. In this work, we propose Diffusion-BBO, a sample-efficient online BBO framework leveraging the conditional diffusion model as the inverse surrogate model. Diffusion-BBO employs a novel acquisition function Uncertainty-aware Exploration (UaE) to propose scores in the objective space for conditional sampling. We theoretically prove that Diffusion-BBO with UaE achieves a near-optimal solution for online BBO. We also empirically demonstrate that Diffusion-BBO with UaE outperforms existing online BBO baselines across 6 scientific discovery tasks.
@@ 1 Introduction
Practical problems in scientific discovery often involve optimizing a black-box objective function that is expensive to evaluate, seen in fields such as superconducting material design (Fannjiang and Listgarten, 2020), DNA sequence optimization (Daniel et al., 2015), and molecular design (Sanchez-Lengeling and Aspuru-Guzik, 2018). How to achieve a near-optimal solution while minimizing function evaluations is thus a major task in black-box optimization (BBO). To improve sample efficiency, prior works in BBO have largely focused on the online setting where the algorithm can iteratively select candidates in the design space and query the black-box function for evaluation (Turner et al., 2021; Zhang et al., 2021; Hebbal et al., 2019; Mockus, 1974). Most existing algorithms belong to the class of forward approaches, including Bayesian optimization (BO) (Kushner, 1964; Mockus, 1974; Wu et al., 2023; Frazier, 2018), bandit algorithms (Agrawal and Goyal, 2012; Karbasi et al., 2023), conditional sampling approaches (Brookes et al., 2019; Gruver et al., 2024; Stanton et al., 2022), in which a surrogate model is built to iteratively approximate and optimize the black-box function.
@@
However, these approaches often fail in scenarios where valid designs, such as valid protein sequences or molecular structures (Kumar and Levine, 2020), are confined in a small subspace. Such optimization problems become challenging as the optimizer needs to eliminate out-of-distribution or invalid designs. Recently, a novel set of methods, termed inverse approaches, have been proposed to address this issue (Kumar and Levine, 2020; Krishnamoorthy et al., 2023; Kim et al., 2024). Specifically, an inverse surrogate model is constructed to learn a mapping from the objective space back to the design space. By leveraging the state-of-the-art generative models, such as diffusion models (Song et al., 2020), these approaches can effectively learn the data distribution in the design space and facilitate optimization within the data manifold (Kong et al., 2024). Promising performance has been achieved in offline BBO settings (Kumar and Levine, 2020; Lu et al., 2023; Wang et al., 2018), assuming an access to a fixed pre-collected dataset.
Despite the advancements of inverse approaches in offline settings, extending existing methods to online BBO remains challenging due to two unresolved issues: (1) the lack of rigorous uncertainty quantification over inverse surrogate models, and (2) the absence of principled methods to leverage the model uncertainty for acquisition function design in the objective space. These limitations hinder the application of inverse approaches for online BBO, where enhancing sample efficiency is crucial as data acquisition is expensive in tasks such as scientific discovery. In this paper, we propose Diffusion-BBO to tackle the above issues. It is a sample-efficient online BBO framework that leverages the conditional diffusion model as the inverse surrogate model. Diffusion-BBO consists of a novel acquisition function Uncertainty-aware Exploration (UaE), which strategically propose the objective scores within the training data distribution for conditional sampling in the design space. We summarize our main contributions as follows:
@@
  • We provide an uncertainty decomposition into epistemic uncertainty and aleatoric uncertainty for conditional diffusion models. We rigorously analyze how uncertainty propagates throughout the denoising process of the conditional diffusion model.
@@
  • We design the acquisition function UaE for Diffusion-BBO based on the quantified uncertainty of conditional diffusion models. Theoretically, we prove that UaE achieves a balance between targeting higher objective values and minimizing epistemic uncertainty, leading to optimal optimization outcomes.
@@
  • We demonstrate that Diffusion-BBO with UaE achieves state-of-the-art online BBO performance across 6 scientific discovery tasks in Design-Bench and molecular discovery.
@@ 2 Related Work
Online Black-box Optimization.
For online Black-box Optimization (BBO), algorithms are designed to iteratively query the oracle function to obtain new training data (Turner et al., 2021; Zhang et al., 2021; Hebbal et al., 2019; Mockus, 1974). In this setting, most existing algorithms belong to the class of forward approaches, including Bayesian optimization (BO) (Kushner, 1964; Mockus, 1974; Wu et al., 2023; Frazier, 2018), bandit algorithms (Agrawal and Goyal, 2012; Karbasi et al., 2023), and conditional sampling approaches (Brookes et al., 2019; Gruver et al., 2024; Stanton et al., 2022). Liu et al. (2024) further integrate LLM capabilities into the BO framework to enable zero-shot warm-starting and enhance surrogate modeling and candidate sampling. Forward approaches build a surrogate model to approximate and optimize the black-box objective function. However, these approaches can struggle with capturing the data manifold in the design space and avoiding out-of-distribution and invalid inputs (Kumar and Levine, 2020). Song et al. (2022); Zhang et al. (2021) proposed Likelihood-free BO using likelihood-free inference to extend BO to a broader class of models and utilities. This approach directly models the acquisition function without separately performing inference with a surrogate model. However, there is a risk that the acquisition function becomes overconfident.
Offline Black-box Optimization.
Recent studies address BBO in the offline setting using pre-collected datasets without actively querying the oracle function. Existing approaches can be categorized into two types. The first is the forward approach, which aims to improve the likelihood of finding a good design with a surrogate model (Yu et al., 2021; Fu and Levine, 2021; Trabucco et al., 2021; Dao et al., 2024). The second is the inverse approach, which models a distribution conditioned on the target function score. This distribution can be estimated through adaptive step-size updates in gradient-based optimization via reinforcement learning (Chemingui et al., 2024) or zero-sum games (Fannjiang and Listgarten, 2020), or autoregressive modeling (Mashkaria et al., 2023). Our work builds on recent progress in inverse approaches, leveraging conditional diffusion models to better learn the data manifold in the design space (Krishnamoorthy et al., 2023; Kong et al., 2024; Li et al., 2024). In contrast to these recent works, we propose a sample-efficient online BBO algorithm with an acquisition function based on conditional diffusion models.
@@
As an emerging class of generative models with strong expressiveness, diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2020) have been successfully deployed across various domains including image generation (Rombach et al., 2022), reinforcement learning (Wang et al., 2022), robotics (Chi et al., 2023), etc. Notably, through the formulation of stochastic differential equations (SDEs), (Song et al., 2020) provides a unified continuous-time score-based framework for distinctive classes of diffusion models. To steer the generation toward high-quality samples with desired properties, it is important to guide the backward data-generation process using task-specific information. Hence, different types of guidance are studied in prior works (Bansal et al., 2023; Nichol et al., 2021; Zhang et al., 2023), including classifier guidance (Dhariwal and Nichol, 2021) where the classifier is trained externally, and classifier-free guidance (Ho and Salimans, 2022), in which the classifier is implicitly specified. In this work, we employ classifier-free guidance to avoid training a separate classifier model, thereby enabling feasible uncertainty quantification in conditional diffusion models.
@@
Uncertainty quantification (UQ) often relies on probabilistic modeling, with Bayesian approximation and ensemble learning being two popular types of approaches. Bayesian Neural Networks (BNNs) (MacKay, 1992; Neal, 2012; Kendall and Gal, 2017; Zhang et al., 2018) employ variational inference to sample model weights from a tractable distribution and estimate uncertainty through sample variance. When training large-scale models, Monte Carlo dropout (Srivastava et al., 2014) offers a cost-effective alternative by approximating BNNs during inference (Gal and Ghahramani, 2016). On the other hand, deep ensembles (Lakshminarayanan et al., 2017) train multiple NNs with different initial weights to gauge uncertainty via model variance. Recent efforts incorporate ensembling techniques in generative models to decompose uncertainty into aleatoric and epistemic components (Valdenegro-Toro and Mori, 2022; Ekmekci and Cetin, 2023). To further improve the scalability of deep ensembles, (Chan et al., 2024) proposed hyper-diffusion to quantify the uncertainty with a single diffusion model. In comparison, we take one step further by utilizing the quantified uncertainty of conditional diffusion models to solve the black-box optimization problem as a downstream task.
3 Preliminaries
3.1 Problem Formulation
@@
Let $f:\mathcal{X}\rightarrow\mathbb{R}$ denote the unknown ground-truth black-box function that evaluates the objective score of any design point $\bm{x}$, with $\mathcal{X}\subseteq\mathbb{R}^{d}$. Our goal is to find the optimal design $\bm{x}^{*}$ that maximizes $f$:
$\bm{x}^{*}\in\operatorname*{argmax}\nolimits_{\bm{x}\in\mathcal{X}}f(\mathbf{\bm{x}}).$ (1)
@@
We are interested in the online BBO setting in which $f$ is expensive to evaluate and the number of evaluations is limited. With a fixed query budget of $K$ and batch size $N$, we iteratively query $f$ with $N$ new inputs in each batch, and update the surrogate model of $f$ based on observed outputs within $K$ iterations. A key concept in online BBO is the acquisition function, which guides the selection of new design points by balancing exploration and exploitation. This function aims to identify high-performing designs, thereby improving the sample efficiency for online BBO.
3.2 Conditional Diffusion Model
Diffusion Models (Sohl-Dickstein et al., 2015; Song et al., 2020) are probabilistic generative models that learn distributions through a denoising process. These models consist of three components: a forward diffusion process that produces a series of noisy samples by adding Gaussian noise, a reverse process to reconstruct the original data samples from the noise, and a sampling procedure to generate new data samples from the learned distribution. Let the original sample be $\mathbf{x}_{0}$ and $t$ be the diffusion step. For conditional diffusion models, the conditioning variable $y$ resides in the same objective space as $f(x)$. It is added to both the forward process as $q\left(\bm{x}_{t}|\bm{x}_{t-1},y\right)$ and the reverse process as $p_{\theta}\left(\bm{x}_{t-1}\mid\bm{x}_{t},y\right),~{}\forall t\in[T]$. The reverse process begins with the standard Gaussian distribution $p(\bm{x}_{T})=\mathcal{N}(\bm{0},\bm{I})$, and denoises $\bm{x}_{t}$ to recover $\bm{x}_{0}$ through the following Markov chain with reverse transitions:
During training, $\Sigma_{\theta}$ is empirically fixed, and $\mu_{\theta}$ is reparametrized by a trainable denoise function $\bm{\epsilon}_{\theta}\left(\bm{x}_{t},t,y\right)$. This function estimates the noise vector $\epsilon$ added to the input $\bm{x}_{t}$, and is trained by minimizing a reweighted version of the evidence lower bound (ELBO):
Notably, the objective function in Equation 2 (Ho et al., 2020) resembles denoising score matching across all time steps $t$. It estimates the gradient of the log probability density of the noisy data (i.e. score function): $\bm{\epsilon}_{\theta}\left(\bm{x}_{t},t,y\right)\approx-\sigma_{t}\nabla_{\bm{x}}\log p(\bm{x}\mid y)$. We further denote the score function as $s_{\theta}(\bm{x}_{t},y,t):=-\bm{\epsilon}_{\theta}\left(\bm{x}_{t},t,y\right)/\sigma_{t}$.
4 Diffusion-BBO
In this section, we present the Diffusion-BBO framework, followed by the details of training its conditional diffusion model in the Bayesian setting.
Problem Setting Surrogate Model Type Acquisition Function Method
Offline BBO Forward Model: $p(y|x,\mathcal{D})$$\times$ Yu et al.; Fu and Levine; Dao et al.
Inverse Model: $p(x|y,\mathcal{D})$$\times$ Chemingui et al.; Kong et al.; Li et al.
Online BBO Forward Model: $p(y|x,\mathcal{D})$ $\alpha(x,\mathcal{D})$ Frazier; Gruver et al.; Song et al.
Inverse Model: $p(x|y,\mathcal{D})$ $\alpha(y,\mathcal{D})$ Our method: Diffusion-BBO
Table 1: Overview of black-box optimization settings, highlighting differences between Diffusion-BBO and existing methods in terms of problem setting, surrogate modeling type, and acquisition function. Note that offline BBO and online BBO are distinct problem settings. We compare Diffusion-BBO only with other online BBO methods. Full discussion of the existing methods is provided in Section 2.
@@ 4 Diffusion-BBO
4.1 The Diffusion-BBO Framework
Diffusion-BBO is an online BBO framework that adopts inverse surrogate models. It leverages the conditional diffusion model to learn the conditional distribution of $p(\bm{x}|y,\mathcal{D})$ with training data $\mathcal{D}$. It supports high-quality conditional sampling given an arbitrary conditioning variable $y$ within the training data distribution. We can design the acquisition function to select $y$ for conditional sampling in the design space. Table 1 compares forward and inverse approaches for solving online and offline BBO problems. The key distinction is that online BBO requires the design of an acquisition function to actively query new data for sequential decision-making. To the best of our knowledge, Diffusion-BBO is the first online BBO approach to propose an acquisition function, $\alpha(y,\mathcal{D})$, for the inverse surrogate model $p(x|y,\mathcal{D})$. In this setting, the objective in Equation 1 becomes:
$\max_{y_{k}\in\mathbb{R}}\sum\nolimits_{k=1}^{K}f(\bm{x}_{k}),~{}~{}\bm{x}_{k}\sim p_{\theta}(\cdot\mid y_{k},\mathcal{D}),~{}~{}\theta\in\Theta.$ (3)
To solve the above optimization problem, we introduce Diffusion-BBO in Algorithm 1. At each iteration $k$, we train a conditional diffusion model and select $y_{k}^{*}$ with the highest acquisition function score $\alpha(y_{k},\mathcal{D})$. In practice, we select $y_{k}$ from a constructed candidate set $\mathcal{Y}_{k}$, where we have $y_{k}=w\cdot\phi_{k}$. The weight $w$ belongs to a fixed set of positive scalars $\mathcal{W}$ and $\phi_{k}$ is the maximum function values being queried in the current training dataset $\mathcal{D}$. Conditioning on $y_{k}^{*}=\arg\max_{y_{k}}{\alpha(y_{k},\mathcal{D})}$, we generate $N$ samples $\{\mathbf{x}_{j}\}_{j=1}^{N}$, where $\bm{x}_{j}\sim p_{\theta}(\mathbf{x}|y_{k}^{*},\mathcal{D})$. By querying the black-box oracle to evaluate each $\bm{x}_{j}$, we obtain the best possible reconstructed value $\phi_{k}$ for the current iteration, and append all queried data pairs $\{\mathbf{x}_{j},f(\mathbf{x}_{j})\}_{j=1}^{N}$ to the training dataset $\mathcal{D}$. The overall Diffusion-BBO framework with the conditional diffusion model is shown in Figure 1.